OpenStudy (anonymous):

Help? Look at the table of values below. a | b ------ 2 8 4 14 6 20 Which equation is represented by the table? A. b = a + 14/2 B. b = 3a + 2 C. b = 4a - 4 D. b = 4a - 2

OpenStudy (e.mccormick):

One of the easiest ways to solve this type is to plug in values and see results. So if you look at the table and the equations, you have b=? Where ? involves a. So put in one of the a values and see if it works. If it does not, you can cross it off. If it does, try another a. If all work, that is the answer.

OpenStudy (anonymous):

Ok, I'll do that!

OpenStudy (anonymous):

@e.mccormick is it c?

OpenStudy (e.mccormick):

Lets see. 2*4=8, 8-4=4... but for a=2, b=8. Nope.

OpenStudy (anonymous):

Ohhh! D then, right?

OpenStudy (e.mccormick):

So c is out. Ah, lets see. I did not do all of the,. Hehe. No, it also has problems. And A has +14/2, which means 7 an odd number, and all these are even. I think you found it by elimination! But the goal is to test and eliminate rather than to guess and eliminate.

OpenStudy (anonymous):

oi how can it not be D? I don't understand? :(

OpenStudy (e.mccormick):

b = 4a - 2 Start with a=2. 4(2)-2=8-2=6. On the chart, it is supposed to be 8 but we got 6.

OpenStudy (anonymous):

Ah I've got it! Massive thank you!

OpenStudy (e.mccormick):

np. Now you get how plugging it in and trying it works. This tactick, while sometimes slow, can solve a large number of questions.